OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x – 4)(2x + 1)(x + 7)?

OpenStudy (ghazi):

just check out the values of X at which the value of function is zero...like at X= 4 the value of function is zero..similarly you can find other values

Parth (parthkohli):

\[ (x - 4)(2x + 1)(x + 7) = 0\]This has 3 zeros. Calculate them by solving these equations:\[x - 4 = 0 \]\[ 2x + 1 = 0\]\[x + 7 = 0 \]

OpenStudy (anonymous):

So you need a positive 4 and negative 7 correct?

OpenStudy (anonymous):

and a negative 1/2

OpenStudy (asnaseer):

yes you are correct :)